Find the corresponding area between z=0 and each of the following:




a) z=0.96



b) z=1.74



c) z=2.18



d) z=2.69



e) z=3.00


1
Expert's answer
2022-03-11T03:20:19-0500

a. "P(Z \leq 0.96)=0.8315" from the z distribution table

"P(Z \leq 0)=0.5000"

The are between is "0.8315-0.5000=0.3315"

b. "P(Z \leq 1.74)=0.9591"

The area becomes "0.9591-0.5000=0.4591"

c. "P(Z \leq 2.18)=0.9854"

The area becomes "=0.9854-0.500=0.4854"

d. "P(Z\leq 2.69)=0.9964"

The area becomes "=0.9964-0.500=0.4964"

e. "P(Z \leq 3.00)=0.9987"

The area becomes "=0.9987-0.500=0.4987"